OpenStudy (anonymous):

Write a letter to a pre-algebra student explaining why algebra problems involve letters. Give at least two examples of expressions involving letters, and explain what the letters used in the examples stand for.

OpenStudy (dangerousjesse):

Remember how you used to use the little boxes in elementary school? \[1+[~~ ]=4\] Well, letters are a lot like those little boxes, only more organized. Can you imagine writing \[6^x t* p+2=x+9*8\] with those little boxes? You'd have to make up all different forms of the box in order for the student to be able to find what [ ] equals. So, now that you're starting to learn about algebra, you need to use letters to help you organize these more complex equations.

OpenStudy (anonymous):

I think I just peed my pants.

OpenStudy (anonymous):

I already did

OpenStudy (anonymous):

Did it feel good?

OpenStudy (anonymous):

nice and warm

OpenStudy (anonymous):

Look sharp ^.^

OpenStudy (dangerousjesse):

Shut it, you're like 12 e.e

OpenStudy (anonymous):

Jesse, I'm 9. FFS get it right :/

OpenStudy (dangerousjesse):

Gah, I always forget. Cut me some slack bruh.

OpenStudy (anonymous):

**Looks Around** Jesus... let's not do this here... I told you I was going to play child support omg...

OpenStudy (anonymous):

Im about 2000% done

OpenStudy (dangerousjesse):

2,000% done= 2,000% cliche

OpenStudy (anonymous):

Sweg af. But to answer this question, say you have .7 on you and your local police officer is trying to cop for $10. Let's also say your mom gave you $x for lunch (that you never used) and you're planning on using that money for a hooker that charges $30 for a full package. Let's create an equation: 10+x=$30

OpenStudy (camerondoherty):

LMAO

OpenStudy (dangerousjesse):

Wait, but what about bus fair? The equation needs to be\[10+x=30+y\]

OpenStudy (anonymous):

True. Stoners don't drive.

OpenStudy (anonymous):

You also forgot the cost of bail from jail after you laced that weed with spice. You're also going to have to figure the probability of picking the jail cell lock before calculating bail in.

OpenStudy (anonymous):

Well, looks like we'd have to convert to Hilbert space. So using the probability of picking jail lock, we can apply the principle of uncertainty and then find the partial derivative of f(x,y). Then we can go ahead and model a contour plot using the function variable of current pocket balance and analyze deltax.

OpenStudy (sedatefrog712):

What does all this have to do with the question

OpenStudy (sedatefrog712):

By talking about stoners bail jail and stuff

OpenStudy (anonymous):

pellet... looks like we'd have to calculate the probability of certainty @ several Cartesian coordinates.

OpenStudy (anonymous):

Well, we can just use the partial of the original probability and apply that over the original derivative

OpenStudy (anonymous):

From all this I think I can actually generate a plausible equation... I'm actually going to try this xD

OpenStudy (anonymous):

Assuming the jail cell was picked (without factoring probability) \[\frac{ \partial f}{ \partial y }(x,y)=\frac{ \Delta f(x)-\Delta x }{ \Delta y + x0 }\]

OpenStudy (anonymous):

Prime*

OpenStudy (anonymous):

Wait that doesn't sound right because we didn't find the limit of the original

OpenStudy (camerondoherty):

Yes so... if it was \(\Large\frac{ \partial f}{ \partial y }(x,y)=\frac{ \Delta f(x)-\Delta x }{ \Delta y + x0 }\) Then it had to go out as: \[\Large \frac{ \Delta \infty \eta \int\limits_{4}^{x} }{ \sum_{d}^{f} }\pm \subseteq \frac{ \prod_{2}^{4} }{ x \rightarrow \lim_{23 \rightarrow 87} }\] Assuming that the bus was ridden and the jailbait was wide open allowing a passage way throught the openings... But if there were no openings how would he subsequently get the bus ticket without paying if he had no money... Im stumped :/

OpenStudy (camerondoherty):

or am i wrong happy?

OpenStudy (anonymous):

What the flutter. Delta, Infinite, nth term and then a false attempt of incomplete integration of no function. Die in a bush Cam.

OpenStudy (camerondoherty):

LOOOOOL

OpenStudy (anonymous):

What you did was basically doing this: sdifhqeuiwtnjgkrquthejwri xD

OpenStudy (camerondoherty):

XD

OpenStudy (camerondoherty):

i got a medal too XD